# Entropy Collapse in Mobile Sensors: The Hidden Risks of Sensor-Based Security

### v10

@@ Abstract
Mobile sensor data has been proposed for security-critical applications such as device pairing, proximity detection, and continuous authentication. However, the foundational premise that these signals provide sufficient entropy remains under-explored. In this work, we systematically analyse the entropy of mobile sensor data using four datasets from multiple application contexts (UCI-HAR, SHL, Relay, and PerilZIS). Using direct computation and estimation, we report entropy values—max, Shannon, collision, and min-entropy—for an exhaustive range of sensor combinations. We demonstrate that the entropy of mobile sensors remains far below what is considered secure by modern standards for security applications, such as for authentication and key generation systems, even when many sensors are combined. In particular, we observe an alarming divergence between average-case Shannon entropy and worst-case min-entropy. Single-sensor mean min-entropy varies between 3.408–4.483 bits despite Shannon entropy being several multiples higher. We further show that correlations between sensor modalities contribute to a $\approx$75% reduction between Shannon and min-entropy. This brings joint min-entropy well below 10 bits in many cases and, in the best case, yielding only $\approx$24 bits when combining 20+ sensor modalities. Our results reveal that adversaries may feasibly predict sensor signals through an exhaustive exploration of the measurement space. Our work also calls into question the widely held assumption that adding more sensors inherently yields higher security. Ultimately, we strongly urge caution when relying on mobile sensor data for security applications.
@@ 1 Introduction
Despite these advances, a crucial assumption remains under-examined: that sensor data is robust against adversaries that aim to forge or reconstruct such data through statistical inference. Many existing proposals rely on heuristic assessments or model classification metrics, e.g. markantonakis2024using , mehrnezhad2015tap , gurulian2018good , shrestha2014drone , truong2014comparing , abuhamad2020sensor , mayrhofer2021adversary , which do not address two basic questions: (1) how much entropy do sensors actually provide and (2) to what extent do multimodal sensor combinations improve entropy? For the latter, multimodal sensing has been proposed as a means to counteract sensor-specific weaknesses shrestha2014drone , markantonakis2024using , truong2014comparing , mehrnezhad2015tap , but the quantitative security benefits of combining multiple sensors have not been rigorously evaluated. The security implications of low-entropy sensor data can be both direct and severe. If such data is used as authentication evidence or to derive secret keys, an adversary may attempt to enumerate the entire space of possible measurement values in order to (re-)derive it. Modern standards provide important quantitative thresholds for what can be considered ‘secure’ for such reasons. For example, NIST SP 800-57 nistsp80057 states that security levels below 80 bits are considered insecure for long-term use, with 112 bits as the current minimum and 128 bits recommended for cryptographic keys. Standards that govern random number generators (RNGs) are stricter still. NIST SP 800-90B turan2018recommendation effectively requires $\approx$1 bit of min-entropy per output bit ($\epsilon\leq 2^{-32}$), while AIS 20/31 bsi2024ais31 mandates over 240 and 250 bits of min-entropy and Shannon entropy respectively for an RNG’s internal state.
@@ 2.1 Related Work
Mobile sensing has been widely explored for security tasks in three broad families: (1) proximity detection and relay-attack mitigation, where co-presence is inferred from shared motion or ambient context halevi2012secure , mehrnezhad2015tap , shepherd2017applicability , gurulian2017effectiveness , markantonakis2024using , gurulian2018good , shrestha2014drone ; (2) device pairing, where two devices derive evidence of co-location to establish a secure association or shared secret mayrhofer2009shake , pan2018universense , xu2021key , mayrhofer2021adversary ; and (3) continuous authentication, which aims to verify users passively during normal interaction riva2012progressive , shi2011senguard , miettinen2014conxsense , li2013unobservable , patel2016continuous , mekruksavanich2021deep , li2018sensor , micallef2015aren . Representative designs include tap-to-pair schemes using accelerometers mehrnezhad2015tap , induced shared vibrations gurulian2018good , and ambient-signal approaches that leverage temperature, humidity, magnetic field, or other environmental cues shrestha2014drone , markantonakis2024using . These mechanisms commonly fuse accelerometer and gyroscope measurements to capture correlated motion as evidence of co-presence mayrhofer2009shake , pan2018universense .
@@
Our key observation is that model-performance metrics are typically decoupled from the entropy of the underlying data sources upon which such models are based. Across the above application areas, effectiveness is usually inferred from discriminative performance, i.e. the system’s performance in classifying illegitimate vs. legitimate samples. A common pipeline—documented across proximity detection, pairing, and continuous authentication—collects data from $N$ participants; derives time- or frequency-domain features (e.g. cross-correlation, spectral energy, and Hamming, Euclidean, and mean absolute distances); learns a decision rule (thresholds or supervised models such as SVMs, Random Forests, Naïve Bayes); and reports FPR/FNR, EER, ROC/AUC, accuracy, or F1. This template appears in numerous systems, including mehrnezhad2015tap , shrestha2014drone , markantonakis2024using , shepherd2017applicability , riva2012progressive , li2013unobservable , miettinen2014conxsense , patel2016continuous , mekruksavanich2021deep , gurulian2018good . For a comprehensive treatment of the model evaluation metrics used by existing mobile sensor-based security systems, the reader is referred to the surveys by Abuhamad et al. abuhamad2020sensor , Alzubaidi and Kalita alzubaidi2016authentication , and Mayrhofer and Sigg mayrhofer2021adversary . Some proposals have attempted to assess the entropy of sensor signals within the context of a security mechanism, but this represents a small minority of the literature. Specifically, T2Pair by Li et al. li2020t2pair reports 32.3–38.5 bits of Shannon entropy, and a refinement by Wu et al. wu2024t2pair reports 51–54 bits. We are aware of no work that applies min-entropy as a worst-case measure of unpredictability turan2018recommendation , bsi2024ais31 , buller2016estimating , alzubaidi2016authentication , xu2021key , mayrhofer2021adversary . Adjacent evidence also highlights practical limits under tight timing constraints. For NFC transactions, EMV specifies $\approx$500 ms for proximity checks. In that setting, Markantonakis et al. markantonakis2024using (building on gurulian2017effectiveness , shepherd2017applicability ) report EERs of 0.179–0.246 across sensor combinations, concluding that commodity sensors are unsuitable for reliable proximity and relay detection without incurring usability or security issues. This points to wider fragility issues with sensor-based security schemes in practice.
@@
Some studies have attempted to assess the entropy bestowed by mobile sensors on a single-sensor basis. Voris et al. voris2011accelerometers investigated accelerometers as true random number generators (TRNGs) on a WISP RFID tag and Nokia N97 phone. The authors find that min-entropy is proportional to the motion applied to the device, with stationary movement having the lowest min-entropy. Intrinsic noise from the sensor circuitry and seismic activity, and the sampling rate of its analog-to-digital converter (ADC), are found to influence entropy generation. Min-entropy values of 3.1–11.4 bits were measured depending on the movement of the accelerometer. Lv et al. lv2020analysis analysed three mobile sensors—a triaxial accelerometer, gyroscope, and magnetometer—on an undisclosed Xiaomi Redmi smartphone. Min-entropy values of 0.593–5.876 are reported, depending on the modality and the entropy estimation method. Krhovják et al. krhovjak2007sources examined image and audio data collected from mobile phone cameras and microphones, respectively, on a Nokia N73 and E-Ten X500 and M700 phones. The work reports Shannon entropies of 2.9 (microphone) and 2.408–5.376 (camera), and min-entropies of 0.5 (microphone) and 0.754–3.928 (camera). Hennebert et al. hennebert2013entropy analysed 10 sensors on two wireless sensor monitors: a TI eZ430-RF2500 and a Zolertia Z1. Min-entropy values of 0–7.85 are given, with motion sensors, e.g. accelerometer and vibration sensors, producing the highest entropy.
@@
Sensors have also been suggested as entropy sources in low-cost RNG designs for mobile devices. Suciu et al. suciu2011unpredictable proposed combining GPS along with accelerometer, gyroscope and orientation sensors on an HTC Google Nexus One, passing NIST SP 800-22 tests rukhin2001nist , but without precise entropy estimates. Wallace et al. wallace2016toward explored a multi-sensor RNG design (accelerometer, gyroscope, microphone, WiFi, GPS, camera) with a non-standard entropy evaluation using 37 Android devices. Unlike dedicated TRNGs (e.g. ring oscillators, Johnson–Nyquist thermal noise, quantum sources hurley2020quantum ), mobile sensors depend heavily on user behaviour and environmental factors, introducing biases and correlations absent in controlled entropy sources. Existing sensor-based mechanisms largely overlook these dynamics, relying principally on heuristic or model evaluation metrics that do not account well for skewed distributions and other biases affecting predictability. In contrast, proposals of new entropy sources and authentication mechanisms commonly rely on worst-case metrics, such as min-entropy as stipulated in NIST SP 800-90B and AIS 20/31 mai2017guessability , uellenbeck2013quantifying , hurley2020quantum , turan2018recommendation , bsi2024ais31 . Our study bridges the gap by systematically evaluating entropy across modalities and datasets.
@@
Voris et al. voris2011accelerometers Acc. $H_{\infty}$ 3.1–11.4 True random number generation
@@
Lv et al. lv2020analysis Acc., Gyr., Mag. $H_{\infty}$ 0.593–5.876 Entropy source analysis on smartphone
Krhovják et al. krhovjak2007sources Cam., Mic. $H_{\{1,\infty\}}$ 0.5–5.376 Randomness sources in mobile devices
Li et al. li2020t2pair Mic., Light $H_{1}$ 32.3–38.5 Zero-interaction device pairing
Wu et al. wu2024t2pair Mic., Light $H_{1}$ 51–54 Zero-interaction device pairing
@@
    Acc.: Accelerometer; Gyr.: Gyroscope; Mag.: Magnetometer; Cam.: Camera; Mic.: Microphone.
@@
In general, while several studies have investigated the entropy of specific sensors, a direct comparison is often difficult due to varying methodologies and contexts. Table 1 provides a structured summary of key prior works, contrasting the sensors analysed, entropy metrics used, and the reported entropy ranges.
@@ 3.2 High-level Methodology
    Data acquisition. We collect large-scale sensor readings from public datasets spanning diverse activities and device usage. Modalities include motion, environmental, and orientation sensors (dataset details are given at the end of this section).
@@
    Relay gurulian2017effectiveness : Approximately 1,500 NFC-based contactless transactions recorded at 100 Hz across several physical locations (e.g. cafés), including accelerometer, gyroscope, and environmental readings taken from emulated payment scenarios.
@@
These datasets provide varied sensor types, user activities, and sampling rates, enabling us to observe how biases and inter-sensor correlations manifest across scenarios. Next, we describe preprocessing and aggregation to form the global distributions used in our entropy analyses.
@@ 4.3 Multi-modal Sensors
Several sensor-based security proposals mehrnezhad2015tap , truong2014comparing , markantonakis2024using , shrestha2014drone , shrestha2018sensor assert that combining multiple sensor modalities bolsters security, on the intuition that an adversary must accurately predict several data streams rather than a single one. An initial, naïve approach may attempt to model this by adding the Shannon entropies from individual sensors, benefitting from the relation $H(X_{1},\ldots,X_{n})=\sum_{i=1}^{n}H(X_{i})$. However, this requires that $X_{i}$ are statistically independent. In reality, mobile sensors often exhibit strong dependencies; for instance, the rotation vector, gravity, and linear acceleration sensors are frequently derived in software from the accelerometer and gyroscope on consumer devices android_motion_sensors . As a result, these modalities cannot be treated as independent variables.
@@
Figure 2 provides a quantitative view of inter-sensor redundancy through Pearson correlation matrices for each dataset. Strong correlations, with $|\rho|\approx 1$, indicate high linear dependence; such redundancy is a major contributor to fragility in the context of pairwise combinations. Employing correlated sensors adds little new unpredictability, which invalidates the simplistic additive model of entropy and can provide diminishing returns. Consequently, while modalities may appear as offering some level of unpredictability individually, the correlations shown in Figure 2 can sharply constrain the overall joint entropy when compared to what would be expected if they were treated as independent variables.

### v12

@@ Abstract
Mobile sensor data has been proposed for security-critical applications such as device pairing, proximity detection, and continuous authentication. However, the foundational premise that these signals provide sufficient entropy remains under-explored. In this work, we systematically analyse the entropy of mobile sensor data using four datasets from multiple application contexts (UCI-HAR, SHL, Relay, and PerilZIS). Using direct computation and estimation, we report entropy values—max, Shannon, collision, and min-entropy—for an exhaustive range of sensor combinations. We demonstrate that the entropy of mobile sensors remains far below what is considered secure by modern standards for security applications, even when many sensors are combined. In particular, we observe an alarming divergence between average-case Shannon entropy and worst-case min-entropy. Single-sensor min-entropy varies between 3.408–4.483 bits despite Shannon entropy being several multiples higher. We also show that redundancies between sensor modalities contribute to a $\approx$75% reduction between Shannon and min-entropy. Indeed, min-entropy plateaus between 8.1–23.9 bits when combining up to 22 modalities, while Shannon entropy can exceed 80 bits. Adding sensors typically increases Shannon entropy but moves min-entropy by only $\approx$1–2 bits per added modality, evidencing entropy collapse under redundancy. Our results reveal that adversaries may feasibly predict sensor signals through an exhaustive exploration of the measurement space. Our work also calls into question the widely held assumption that adding more sensors inherently yields higher security. Ultimately, we strongly urge caution when relying on mobile sensor data for security applications.
@@ 1 Introduction
Despite these advances, a crucial assumption remains under-examined: that sensor data is robust against adversaries aiming to forge or reconstruct such data through statistical inference. Existing proposals rely heavily on heuristic assessments or model classification metrics, e.g. markantonakis2024using , mehrnezhad2015tap , gurulian2018good , shrestha2014drone , truong2014comparing , abuhamad2020sensor , mayrhofer2021adversary , which do not address two basic questions: (1) how much entropy do sensors actually provide and (2) to what extent do multimodal sensor combinations improve entropy? For the latter, multimodal sensing has been proposed as a means to counteract sensor-specific weaknesses shrestha2014drone , markantonakis2024using , truong2014comparing , mehrnezhad2015tap , but the quantitative security benefits of combining multiple sensors have not been rigorously evaluated. The security implications of low-entropy sensor data can be both direct and severe. If such data is used as authentication evidence, or to derive secret keys, an adversary may attempt to enumerate the entire space of possible sensor measurement values in order to recover authentication or key material. Modern standards provide important quantitative thresholds for what can be considered ‘secure’ for such reasons. For example, NIST SP 800-57 nistsp80057 states that security levels below 80 bits are considered insecure for long-term use, with 112 bits as the current minimum and 128 bits recommended for cryptographic keys. Standards that govern random number generators (RNGs) are stricter still. NIST SP 800-90B turan2018recommendation effectively requires $\approx$1 bit of min-entropy per output bit ($\epsilon\leq 2^{-32}$), while AIS 20/31 bsi2024ais31 mandates over 240 and 250 bits of min-entropy and Shannon entropy respectively for an RNG’s internal state.
@@ 2.1 Related Work
Mobile sensing has been widely explored for security tasks in three broad families: (1) proximity detection and relay-attack mitigation, where co-presence is inferred from shared motion or ambient context halevi2012secure , mehrnezhad2015tap , shepherd2017applicability , gurulian2017effectiveness , markantonakis2024using , gurulian2018good , shrestha2014drone ; (2) device pairing, where two devices derive evidence of co-location to establish a secure association or shared secret mayrhofer2009shake , pan2018universense , xu2021key , mayrhofer2021adversary ; and (3) continuous authentication, which aims to verify users passively during normal interaction riva2012progressive , shi2011senguard , miettinen2014conxsense , li2013unobservable , patel2016continuous , mekruksavanich2021deep , li2018sensor , micallef2015aren . Representative designs include tap-to-pair schemes using accelerometers mehrnezhad2015tap , induced shared vibrations gurulian2018good , and ambient-signal approaches that leverage temperature, humidity, magnetic field, or other environmental cues shrestha2014drone , markantonakis2024using . Proposals commonly combine sensor modalities to gather additional evidence of co-presence; for example, using accelerometer and gyroscope measurements to capture shared device motion in addition to environmental readings mayrhofer2009shake , pan2018universense , alzubaidi2016authentication , xu2021key .
@@
A common pipeline—documented across proximity detection, pairing, and continuous authentication—collects data from $N$ participants; derives time- or frequency-domain features (e.g. cross-correlation, spectral energy, and Hamming, Euclidean, and mean absolute distances); learns a decision rule (thresholds or supervised models such as SVMs, Random Forests, Naïve Bayes); and reports FPR/FNR, EER, ROC/AUC, accuracy, or F1. This template appears in numerous systems, including mehrnezhad2015tap , shrestha2014drone , markantonakis2024using , shepherd2017applicability , riva2012progressive , li2013unobservable , miettinen2014conxsense , patel2016continuous , mekruksavanich2021deep , gurulian2018good . Across the above application areas, effectiveness is usually inferred from discriminative performance, i.e. the system’s performance in classifying illegitimate vs. legitimate samples. (For a comprehensive treatment of the model evaluation metrics used by existing mobile sensor-based security systems, the reader is referred to the surveys by Abuhamad et al. abuhamad2020sensor , Alzubaidi and Kalita alzubaidi2016authentication , and Mayrhofer and Sigg mayrhofer2021adversary ). Our key observation is that model-performance metrics are typically decoupled from the entropy of the underlying data sources upon which such models are based. Some proposals have attempted to assess the entropy of sensor signals within the context of a security mechanism, but this represents a small minority of the literature. Specifically, T2Pair by Li et al. li2020t2pair reports 32.3–38.5 bits of Shannon entropy, and a refinement by Wu et al. wu2024t2pair reports 51–54 bits. We are aware of no work that applies min-entropy as a worst-case measure of unpredictability turan2018recommendation , bsi2024ais31 , buller2016estimating , alzubaidi2016authentication , xu2021key , mayrhofer2021adversary . Adjacent evidence also highlights practical limits under tight timing constraints. For NFC transactions, EMV specifies $\approx$500 ms for proximity checks. In that setting, Markantonakis et al. markantonakis2024using (building on gurulian2017effectiveness , shepherd2017applicability ) report EERs of 0.179–0.246 across sensor combinations, concluding that commodity sensors are unsuitable for reliable proximity and relay detection without incurring usability or security issues. This points to wider fragility issues with sensor-based security schemes in practice.
@@
Some studies have attempted to measure the entropy of single mobile sensors. Voris et al. voris2011accelerometers evaluated accelerometers as true RNGs (TRNGs) on a WISP RFID tag and Nokia N97 phone. The authors find that min-entropy is proportional to the motion applied to the device, with stationary movement having the lowest min-entropy. Intrinsic noise from the sensor circuitry and seismic activity, and the sampling rate of its analog-to-digital converter (ADC), are found to influence entropy generation. Min-entropy values of 3.1–11.4 bits were measured depending on the movement of the accelerometer. Lv et al. lv2020analysis analysed three mobile sensors—a triaxial accelerometer, gyroscope, and magnetometer—on an undisclosed Xiaomi Redmi smartphone. Min-entropy values of 0.593–5.876 are reported, depending on the modality and the entropy estimation method. Krhovják et al. krhovjak2007sources examined image and audio data collected from mobile phone cameras and microphones, respectively, on a Nokia N73 and E-Ten X500 and M700 phones. The work reports Shannon entropies of 2.9 (microphone) and 2.408–5.376 (camera), and min-entropies of 0.5 (microphone) and 0.754–3.928 (camera). Hennebert et al. hennebert2013entropy analysed 10 sensors on two wireless sensor monitors: a TI eZ430-RF2500 and a Zolertia Z1. Min-entropy values of 0–7.85 are given, with motion sensors, e.g. accelerometer and vibration sensors, producing the highest entropy.
@@
Sensors have also been suggested as entropy sources within low-cost RNG designs for mobile devices. Suciu et al. suciu2011unpredictable proposed combining GPS along with accelerometer, gyroscope and orientation sensors on an HTC Google Nexus One, passing NIST SP 800-22 tests rukhin2001nist , but without precise entropy estimates. Wallace et al. wallace2016toward explored a multi-sensor RNG design (accelerometer, gyroscope, microphone, WiFi, GPS, camera) with a non-standard entropy evaluation using 37 Android devices. Unlike dedicated TRNGs (e.g. ring oscillators, Johnson–Nyquist thermal noise, quantum sources hurley2020quantum ), mobile sensors depend heavily on user behaviour and environmental factors, introducing biases and correlations absent in controlled entropy sources. Existing sensor-based mechanisms largely overlook these dynamics, relying principally on heuristic or model evaluation metrics that do not account well for skewed distributions and other biases affecting predictability. In contrast, proposals of new entropy sources and authentication mechanisms commonly rely on worst-case metrics, such as min-entropy as stipulated in NIST SP 800-90B and AIS 20/31 mai2017guessability , uellenbeck2013quantifying , hurley2020quantum , turan2018recommendation , bsi2024ais31 . While several studies have investigated the entropy of specific sensors, a direct comparison is often difficult due to varying methodologies and contexts. Table 1 provides a structured summary of key prior works, contrasting the sensors analysed, entropy metrics used, and the reported entropy ranges. Our study bridges the gap by systematically evaluating entropy across modalities and datasets.
@@
Voris et al. voris2011accelerometers Acc $H_{\infty}$ 3.1–11.4 True random number generation
@@
Lv et al. lv2020analysis Acc, Gyro, Mag $H_{\infty}$ 0.593–5.876 Entropy source analysis on smartphone
Krhovják et al. krhovjak2007sources Cam, Mic $H_{\{1,\infty\}}$ 0.5–5.376 Randomness sources in mobile devices
Li et al. li2020t2pair Mic, Light $H_{1}$ 32.3–38.5 Zero-interaction device pairing
Wu et al. wu2024t2pair Mic, Light $H_{1}$ 51–54 Zero-interaction device pairing
@@
    Acc: Accelerometer; Gyro: Gyroscope; Mag: Magnetometer; Cam: Camera; Mic: Microphone.
@@ 3.2 High-level Methodology
    Data acquisition. We collect large-scale sensor readings from public datasets spanning diverse activities and device usage. Modalities include motion, environmental, and orientation sensors (details are given at the end of this section).
@@
    Relay markantonakis2024using : Approximately 1,500 NFC-based contactless transactions recorded at 100 Hz across several physical locations (e.g. cafés), including accelerometer, gyroscope, and environmental readings taken from emulated payment scenarios.
@@
These datasets provide varied sensor types, user activities, and sampling rates, enabling us to observe how biases and inter-sensor correlations manifest across scenarios. We summarise the datasets in Table 2.
Table 2: Summary of datasets.
Dataset Devices Sensorsa Freq. Contextb
UCI-HAR anguita2013public Samsung Galaxy S2 Acc, Gyro 50Hz ADL; 30 users
SHL gjoreski2018university , wang2019enabling Huawei Mate 9 Acc, Gyro, Mag, Grav, LinAcc, RotVec, Light, Temp, Pres, Alt 100Hz HAR; transport; 3 users; 59 hrs
Relay markantonakis2024using Google Nexus 5, 9; Samsung Galaxy S4, SGS5 Acc, Gyro, Mag, RotVec, Grav, LinAcc, Light 100Hz NFC transactions; 252 users
PerilZIS fomichev2019perils TI SensorTag; Samsung Galaxy S6, Gear Acc, Gyro, Mag, Light, Temp, Pres 10Hz ZIS; 4 users; 8 devices; 4250 hrs
  • a
    Acc: Accelerometer; Gyro: Gyroscope; Mag: Magnetometer; Grav: Gravity; LinAcc: Linear acceleration; RotVec: Rotation vector; Temp: Temperature; Pres: Pressure; Alt: Altitude.
  • b
    ADL: Activities of daily living; HAR: Human activity recognition; NFC: Near-field communication; ZIS: Zero-interaction security.
@@ 4.3 Multi-modal Sensors
Several sensor-based proposals mehrnezhad2015tap , truong2014comparing , markantonakis2024using , shrestha2014drone , shrestha2018sensor assert that combining multiple sensor modalities bolsters security, on the intuition that an adversary must accurately predict several data streams rather than a single one. An initial, naïve approach may attempt to model this by adding the Shannon entropies from individual sensors, benefitting from the relation $H(X_{1},\ldots,X_{n})=\sum_{i=1}^{n}H(X_{i})$. However, this requires that $X_{i}$ are statistically independent. In reality, mobile sensors often exhibit strong dependencies; for instance, the rotation vector, gravity, and linear acceleration sensors are frequently derived in software from the accelerometer and gyroscope on consumer devices android_motion_sensors . As a result, these modalities cannot be treated as independent variables.
@@
Figure 2 provides a quantitative view of inter-sensor redundancy through Pearson correlation matrices for each dataset. Strong correlations, with $|\rho|\approx 1$, indicate high linear dependence; such redundancy is a major contributor to fragility in the context of pairwise combinations. Employing correlated sensors adds little new unpredictability, which invalidates the simplistic additive model of entropy and can provide diminishing returns. Figure A.1 (Appendix A) presents mutual information heatmaps. These results reveal dependencies between sensor pairs, confirming that redundancy extends beyond what Pearson correlation alone suggests. Consequently, while modalities may appear as offering some level of unpredictability individually, the results suggest that overall joint entropy can be sharply constrained compared to what would be expected if they were treated as independent variables.